12. Two APs have the same common difference. The difference between their 100th term is 100, what is the difference between their 1000th terms?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let, the first term of two APs be a1 and a2 respectively And the common difference of these APs be d. For the first A.P.,we know, an = a+(n−1)d Therefore, a100 = a1+(100−1)d = a1 + 99d a1000 = a1+(1000−1)d a1000 = a1+999d For second A.P., we know, an = a+(n−1)d Therefore, a100 = a2+(100−1)d = a2+99d a1000 = a2+(1000−1)d = a2+999d Given that, difference between 100th term of the two APs = 100 Therefore, (a1+99d) − (a2+99d) = 100 a1−a2 = 100……………………………………………………………….. (i) Difference between 1000th terms of the two APs (a1+999d) − (a2+999d) = a1−a2 From equation (i), This difference, a1−a2 = 100 Hence, the difference between 1000th terms of the two A.P. will be 100.